Aries:

The radius of a circular sign is 12 inches. Equal parts of the sign are painted red and yellow. How many square inches are painted in each color? Use 3.14 for pi.

Renne:

To determine how many square inches are painted in each color on a circular sign with a radius of 12 inches, we first need to find the area of the entire sign. The formula for the area of a circle is A = πr^2, where A is the area and r is the radius. Plugging in the given radius of 12 inches and using 3.14 for pi, we get: A = 3.14 x 12^2 A = 3.14 x 144 A = 452.16 square inches Since the sign is divided into equal parts of red and yellow, we can find the area of each color by dividing the total area by 2. Area of red = 452.16 ÷ 2 = 226.08 square inches Area of yellow = 452.16 ÷ 2 = 226.08 square inches Therefore, equal parts of the 12-inch circular sign are painted in red and yellow, with each color covering an area of 226.08 square inches.

Aries:

What is that in inches?

Renne:

The information is stated within.